OpenStudy (migitmack):

One endpoint of a segment is (12,-8). The midpoint is (3,18). Find the coordinates of the other endpoint.

OpenStudy (anonymous):

Could use the formula, but it's easy enough to just reason through it: What number is the same distance from 3 as 12, and what number is the same distance from 18 as -8?

OpenStudy (migitmack):

Can U just tell me I am kinda lazy -_-

OpenStudy (anonymous):

Nope. Stop being lazy. I will draw a picture for you though.

OpenStudy (anonymous):



OpenStudy (anonymous):

Your other option is to use the formula: \[M(x,y)=\left(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\right).\]

OpenStudy (migitmack):

1 sec

OpenStudy (cathyangs):

@Migitmack , @CliffSedge is explaining it quite well. Openstudy is supposed to help you study, it's not an answer-generator.

OpenStudy (migitmack):

Whats M

OpenStudy (anonymous):

M=midpoint.

OpenStudy (migitmack):

I don't understand

OpenStudy (migitmack):

i did the formula

OpenStudy (migitmack):

but I where is the M

OpenStudy (anonymous):

You have the coordinates for the midpoint and one other point, so rearrange the formula to solve for what you don't know. (If you want to use the formula.. I still think just doing some elementary subtraction is more straight-forward) 

OpenStudy (migitmack):

I hoy (7.5, -72)?!!!!!?! I think I did something wrong

OpenStudy (anonymous):

Just look at the first picture I drew and ask yourself these questions: 1. 3 is halfway between 12 and what other number? That's your x. 2. 18 is halfway between -8 and what other number? That's your y.

OpenStudy (migitmack):

Is X 9 or 7?

OpenStudy (anonymous):

The formula would look like this: \[M(3,18)=(\frac{12+x}{2},\frac{-8+y}{2}).\] Handle the x and y equations separately: \[3=\frac{12+x}{2}, \space 18=\frac{-8+y}{2}\]

OpenStudy (anonymous):

The distance from 3 to 12 is 9, so what number is also 9 units away from 3, but in the other direction? Use the same reasoning for y.

OpenStudy (migitmack):

For Y its going to be 5?

OpenStudy (migitmack):

@CliffSedge

OpenStudy (migitmack):

@Cliffsedge yo is it 5 for y?

OpenStudy (anonymous):

Why do you think 5 for y? How did you get that number?

OpenStudy (anonymous):

I'm guessing you're finding the midpoint of (3,18) and (12,-8) but (3,18) *is* the midpoint.